two boats are going the same speed on beautiful Clear Creek in Colorado. One is heading upstream and travels 60 miles in 3 hours. The other boat is going downstream and is able to travel 132 miles in 3 hours. What is the speed of the water current and the speed of the boats
Answer by Alan3354(69443)

two boats are going the same speed on beautiful Clear Creek in Colorado. One is heading upstream and travels 60 miles in 3 hours. The other boat is going downstream and is able to travel 132 miles in 3 hours. What is the speed of the water current and the speed of the boats
---------
b = speed of the boats WRT the water.
c = current
---
b - c = 60/3 = 20 mi/hr
b + c = 132/3 = 44 mi/hr
==========
etc.